The length of a rectangle is 5 more than the width. The area is 500 square inches. Find the length and width of the rectangle.

Answer by mananth(16946)


The length of a rectangle is 5 more than the width.
let the width be x
the length = (x+5)
x(x+5) =500
x^2+5x=500
x^2+5x-500=0
x^2+25z-20x-500=0
x(x+25)-20(x+25)=0
(x+25)(x-20)=0
x=20 the positive value
20 and 25 in are the dimensions
The area is 500 square inches.